# Protecting Deep Learning Model Copyrights with Adversarial Example-Free Reuse Detection

### v1

@@ III-B Neuron Functionality Analysis-based Reuse Detector
Based on the finding in our preliminary study, we propose NFARD, a neuron functionality analysis-based reuse detector. We depict the detection procedure in Algorithm 1, which consists of three core components: test suite selection, distance metrics, and decision criteria. We first concentrate on dealing with homogeneous cases in white-box and black-box settings and present the linear transformation method for heterogeneous cases later.
@@ III-B1 Test suite selection
NFARD uses a subset of the victim model’s training data to extract neuron vectors to compute the neuron functionality similarity between different neural networks. Note that NFARD does not require adding any perturbations to the test samples. To induce aligning behavior between the victim and surrogate models, we construct the test suite using samples that elicit the least confident predictions from the victim model, where the degree of confidence is measured by the entropy of the predicted probabilities.
Input: victim model $f_{v}$, suspect model $f_{s}$, test suite size $n$, dataset $D$, reference models $\{f_{r,i}\}_{i=1}^{R}$
Output: decision on if $f_{s}$ is a surrogate of $f_{v}$
@@ III-B2 Distance metrics
In the white-box setting, the detector has access to the neuron vectors of all layers, so we can directly adopt $dist_{eu}^{(k)}$ and $dist_{ac}^{(k)}$ as the distance metrics. The parameter $k$ specifies the target layer from which we extract the neuron vectors, and the choice of target layer depends on neural network architecture. In practice, we usually choose a relatively shallow layer as the target layer.
In the black-box setting, neither $dist_{eu}^{(k)}$ nor $dist_{ac}^{(k)}$ is directly applicable as we have the probability predictions (by applying softmax function) rather than the neuron vectors of the last layer. It is impossible to recover the output (neuron vector) of the last layer from the probability predictions since the softmax function is not bijective. We propose to take the logarithm of the probability to compute a logit approximation. The differences in the computation of distance metrics in the black-box setting are shown as follows:
$dist_{eu}^{\prime}(f_{v},f_{s},X)$ $=\dfrac{1}{m}\sum\limits_{i=1}^{m}\lVert\tilde{\bm{\mathrm{nv}}_{v}}(L,i)-\tilde{\bm{\mathrm{nv}}_{s}}(L,i)\rVert_{2},$
$dist_{ac}^{\prime}(f_{v},f_{s},X)$ $=\dfrac{1}{m}\sum\limits_{i=1}^{m}dist_{cos}(\tilde{\bm{\mathrm{nv}}_{v}}(L,i),\tilde{\bm{\mathrm{nv}}_{s}}(L,i)),$
@@
$\tilde{\bm{\mathrm{nv}}}(L,i)=(\log f(\bm{x}_{1})_{i},\log f(\bm{x}_{2})_{i},\cdots,\log f(\bm{x}_{n})_{i})^{\intercal}.$
@@ III-B3 Decision criteria
NFARD requires a set of reference models to make the final decision. For each distance metric $dist_{i}$, we define its decision function as:
$d_{i}(x,y_{1},\cdots,y_{s})=\tilde{z}-\alpha\cdot\textrm{IQR}-x,$
where $x$ is the distance between $f_{v}$ and $f_{s}$, $y_{i}$ is the distance between $f_{v}$ and $f_{r,i}$, $\tilde{z}$ and IQR are the median and the interquartile range of $\{x,y_{1},\cdots,y_{s}\}$, and $\alpha$ is a pre-fixed parameter. Conventionally, data points below $\tilde{z}-1.5\ \textrm{IQR}$ are regarded as outliers (known as the IQR rule). In our method, we replace the coefficient 1.5 with a user-specified parameter to control the trade-off between sensitivity and specificity. The higher the decision function $d_{i}$ value, the greater the deviation of the suspect model from the reference models, indicating a higher likelihood of it being a surrogate model.
@@
$decision(x,y_{1},\cdots,y_{s})=\sum_{i=1}^{p}w_{i}\cdot d_{i}(x,y_{1},\cdots,y_{s}),$
where $w_{i}$ is the weight of the distance metric $dist_{i}$. If the weighted sum is greater than 0, NFARD classifies the suspect model as a true surrogate model. Otherwise, the suspect model is considered an independently trained model.
@@ III-C Handling Heterogeneous Reuse Cases
Without loss of generality, suppose we have two (approximated) neuron matrices $\bm{H}_{1}\in\mathbb{R}^{n\times a}$, $\bm{H}_{2}\in\mathbb{R}^{n\times b}$ and $a\geq b$, where the (approximated) neuron matrix $\bm{H}$ is defined as:
$\bm{H}=\begin{pmatrix}\bm{\mathrm{nv}}(k,1)&\bm{\mathrm{nv}}(k,2)&\cdots&\bm{\mathrm{nv}}(k,m_{k})\end{pmatrix}.$
The main obstacle here is that we cannot directly compare the neuron vectors of $\bm{H}_{1}$ and $\bm{H}_{2}$ because there is no apparent functionality correspondence between the neurons of the two models with different structures. To tackle this problem, we construct a linear transformation $\bm{P}\in\mathbb{R}^{b\times a}$ to project neuron vectors in $\bm{H}_{1}$ into a space of the same dimension as $\bm{H}_{2}$ so that the metrics we designed above become applicable to $\bm{H}_{1}\bm{P}^{\intercal}$ and $\bm{H}_{2}$. The linear transformation $\bm{P}$ should preserve the difference among independently trained models in terms of neuron functionality, as well as the similarity between victim models and surrogate models, to enable reuse relation detection between heterogeneous models.
@@
where $\lVert\cdot\rVert_{F}$ is the Frobenius norm. Intuitively, neuron vectors of reference models are largely linearly independent of victim model’s neuron vectors as they are independently trained, so the distances between victim models and reference models will still remain relatively large after the linear transformation.
@@
$\bm{p}_{i}=\operatorname*{argmin}_{\bm{w}_{i}}\lVert\bm{H}_{1}\bm{w}_{i}-\bm{\mathrm{nv}}_{2}(k,i)\rVert_{2},$
@@ IV Benchmark
In order to evaluate our approach comprehensively, we create a benchmark named Reuse Zoo consisting of 250 models for model reuse detection. This benchmark contains 4 pre-trained models, 126 surrogate models, and 120 reference models with four popular classification model architectures, including ResNet-18 [35], ResNet-50 [35], MobileNet V2 [36], and GoogLeNet [37].
@@
Knowledge distillation and model extraction are typically used to train smaller surrogates with different architectures from the source model. Since the capacity of ResNet-50 is significantly larger than that of the other three architectures, we train three knowledge distillation surrogates and three model extraction surrogates (with ResNet-18, MobileNet V2, and GoogLeNet architectures) for each ResNet-50 transfer learning model. We use the vanilla method [33] for knowledge distillation and the Knockoff method [9] for model extraction. These models constitute the knowledge distillation surrogate set with 9 models and the model extraction surrogate set with 9 models. In total, we obtain a total of 126 surrogate models using these seven model reuse techniques.
We construct 120 reference models in the following way. We independently train ten models from scratch using different weight initialization for each model architecture (four architectures) and each dataset (three datasets). All surrogate and reference models are well-trained and have qualified performance. The complete list of models in the Reuse Zoo benchmark is shown in Table I, with several examples.
This benchmark covers four popular model architectures, ranging from small networks like MobileNet V2 to large networks like ResNet-50, and various reuse methods, notably including heterogeneous reuse methods that can change the model structure. We note that the authors of ModelDiff [25] have proposed a similar DNN reuse detection benchmark referred to as ModelReuse. Compared with ModelReuse, Reuse Zoo (1) is constructed with three commonly used classification datasets while ModelReuse uses two fine-grained image classification (FGIC) datasets that are less representative; (2) includes two more model architectures than ModelReuse, namely ResNet-50 and GoogLeNet, where ResNet-50 is a significantly larger model; (3) covers more reuse techniques, such as fine-tuning and retraining on different layers, which are absent from ModelReuse; (4) is larger, more balanced (126 surrogate models and 120 reference models vs. 84 and 28), and of higher quality (5% accuracy gap between homogeneous models vs. 25% in ModelReuse).
@@
Reuse type Detected Decision $dist_{eu}^{\prime}$ $dist_{ac}^{\prime}$ Detected Decision $dist_{eu}^{(k)}$ $dist_{ac}^{(k)}$
@@
V Empirical Evaluation
Parameter setting. We implement NFARD using PyTorch [42] and release its source code along with the benchmark. [^4] https://figshare.com/s/691d5af98dc715cf7009 for anonymous review NFARD has four parameters: the test suite size $n$, the user-specified parameter $\alpha$ for decision criteria, the distance metrics weights $\{w_{i}\}$, and the choice of target layer under the white-box setting. If not specified below, we use a test suite size of $n=1000$, and we empirically set the weights of Euclidean and cosine distances as 1 and 120, respectively. $\alpha$ is set to 0.85 for the black-box case and 3.5 for the white-box case. The target layer is set to the layer at about 25% depth. We will discuss the choice of $\alpha$ and target layer later.
V-A RQ1: Performance of NFARD
V-A1 Benchmark evaluation result
@@ V-B RQ2: Application Scope
TABLE V: Qualitative comparison of the application scope of NFARD with adversarial example-based methods.
@@
TABLE VI: Test suite generation time (in seconds) of reuse detectors. The shortest generation time is highlighted.
@@
We have shown that NFARD is capable of detecting more reuse types than other DNN reuse detection methods by employing the linear transformation method. To answer the second research question, we further compare the application scope of NFARD with other methods in the following aspects: model access, scalability, and label information. We provide a brief summary in Table V and describe them in detail.
@@ V-B2 Scalability
Another important factor that affects the application scope is scalability, which is mainly limited by the efficiency of test suite generation as the metric computation and decision-making process usually takes no more than a few seconds. We report the average test suite generation time for different model architectures of NFARD and other adversarial example-based methods (with $n=1000$) in Table VI. The test suite generation process for NFARD is the same in black-box and white-box cases, while DeepJudge takes two different procedures for test suite generation in each case. NFARD is $2\sim 99$ times faster than the baseline reuse detectors, and its efficiency is also less influenced by the complexity of the model. Since NFARD directly selects samples from the training dataset, the test suite generation time stays almost the same as the number of test cases grows. On the contrary, the test suite generation time of adversarial example-based reuse detectors grows linearly with the number of test cases.
@@ V-C RQ3: Influencing Factors
Next, we explore how different configurations affect the performance of NFARD. First, we investigate the effectiveness of logarithm approximation in the black-box case by an ablation experiment. Then, we test the performance with different parameters, including decision parameter $\alpha$, target layer choice, and test suite size $n$.
V-C1 Logarithm approximation
TABLE VII: Evaluation results of NFARD without taking the logarithm under the black-box setting.
Reuse type Detected Decision $dist_{eu}^{\prime}$ $dist_{ac}^{\prime}$
@@ V-C1 Logarithm approximation
We report the evaluation results of NFARD without taking the logarithm in the black-box setting in Table VII. Clearly, there is a significant drop in performance without the logarithm approximation, especially for knowledge distillation and model extraction reuse cases. The decision value also decreases for most reuse cases, suggesting that the logarithm approximation method does amplify the difference between independently trained models and surrogate models in the black-box case.

### v2

@@ III-B Neuron Functionality Analysis-based Reuse Detector
Based on the finding in our preliminary study, we propose NFARD, a neuron functionality analysis-based reuse detector. Algorithm 1 demonstrates its procedure, which includes three core components: test suite selection, distance metrics, and decision criteria. We first concentrate on dealing with homogeneous cases in white-box and black-box settings and present the linear transformation method for heterogeneous cases later.
@@ III-B1 Test suite selection
NFARD uses a subset of the victim model’s training data to extract neuron vectors to compute the neuron functionality similarity between different neural networks. Note that NFARD does not require adding any perturbations to the test samples. To induce aligning behavior between the victim and surrogate models, we construct the test suite using samples that elicit the least confident predictions from the victim model. The degree of confidence is measured by the entropy of the predicted probabilities.
Input: victim model $f_{v}(\,\cdot\,;\theta_{v})$, suspect model $f_{s}(\,\cdot\,;\theta_{s})$, size $n$, dataset $D$, reference models $\{f_{r_{i}}(\,\cdot\,;\theta_{r_{i}})\}_{i=1}^{R}$
Output: decision on if $f_{s}(\,\cdot\,;\theta_{s})$ is a surrogate of $f_{v}(\,\cdot\,;\theta_{v})$
@@ III-B2 Distance metrics
In the white-box setting, the detector has access to the neuron vectors of all layers, so we can directly adopt $\mathrm{ED}^{(k)}$ and $\mathrm{CD}^{(k)}$ as the distance metrics. The parameter $k$ specifies the target layer from which we extract the neuron vectors, and the choice of target layer depends on neural network architecture. In practice, we usually choose a relatively shallow layer as the target layer.
In the black-box setting, neither $\mathrm{ED}^{(k)}$ nor $\mathrm{CD}^{(k)}$ is directly applicable as we have the probability predictions (by applying softmax function) rather than the neuron vectors of the last layer. It is impossible to recover the output (neuron vector) of the last layer from the probability predictions since the softmax function is not bijective. We propose to take the logarithm of the probability to compute a logit approximation. The differences in the computation of distance metrics in the black-box setting are shown as follows:
$\mathrm{BED}(f_{v},f_{s})$ $=\dfrac{1}{m}\sum\limits_{i=1}^{m}\lVert\tilde{\bm{\mathrm{nv}}_{v}}(L,i)-\tilde{\bm{\mathrm{nv}}_{s}}(L,i)\rVert_{2},$ (4)
$\mathrm{BCD}(f_{v},f_{s})$ $=\dfrac{1}{m}\sum\limits_{i=1}^{m}(1-\dfrac{\tilde{\bm{\mathrm{nv}}_{v}}(L,i)^{\intercal}\tilde{\bm{\mathrm{nv}}_{s}}(L,i)}{\lVert\tilde{\bm{\mathrm{nv}}_{v}}(L,i)\rVert_{2}\lVert\tilde{\bm{\mathrm{nv}}_{s}}(L,i)\rVert_{2}})$ (5)
@@
$\tilde{\bm{\mathrm{nv}}}(L,i)=(\log f(\bm{x}_{1})_{i},\log f(\bm{x}_{2})_{i},\cdots,\log f(\bm{x}_{n})_{i})^{\intercal}.$ (6)
@@ III-B3 Decision criteria
NFARD requires a set of reference models to make the final decision. For a given distance metric $D_{i}$, the goal is to determine whether the distance between the victim model and the suspect model, $D_{i}(f_{v},f_{s})$, is significantly smaller than the distances between the victim model and the reference models, $\{D_{i}(f_{v},f_{r_{j}})\}_{j=1}^{R}$. To achieve this, we design a threshold function based on the interquartile range (IQR) rule, which identifies data points below the first quartile minus 1.5 times the IQR as outliers. To balance specificity and sensitivity, we generalize this rule by introducing a sensitivity parameter $\alpha$, replacing the coefficient 1.5. The threshold function is defined as the gap between the adjusted IQR lower bound and the distance metric value:
$\tau_{i}(f_{v},f_{s},\{f_{r_{j}}\}_{j=1}^{R})=M_{i}-\alpha\cdot\mathrm{IQR}_{i}-D_{i}(f_{v},f_{s}),$ (7)
where $M_{i}$ is the median of the distances $\{D_{i}(f_{v},f_{s})\}\cup\{D_{i}(f_{v},f_{r_{j}})\}_{j=1}^{R}$, and $\mathrm{IQR}_{i}$ is their interquartile range. The user-defined parameter $\alpha$ controls the trade-off between specificity and sensitivity: higher $\alpha$ values make the criterion stricter, favoring specificity over sensitivity. A higher value of the threshold function $\tau_{i}$ indicates a greater deviation of the suspect model from the reference models, thus a higher likelihood of it being a surrogate model.
@@
$\Psi(f_{v},f_{s},\{f_{r_{j}}\}_{j=1}^{R})=\sum_{i=1}^{p}w_{i}\cdot\tau_{i}(f_{v},f_{s},\{f_{r_{j}}\}_{j=1}^{R}),$ (8)
referred to as the decision value, where $w_{i}$ is the weight of the distance metric $D_{i}$ and there are $p$ distance metrics in total. If the weighted sum is positive, NFARD identifies the suspect model as a surrogate model; otherwise, it is classified as an independently trained model.
@@ III-C Handling Heterogeneous Reuse Cases
Without loss of generality, suppose we have two sets of neuron vectors from two models, $\{\bm{\mathrm{nv}}_{1}(k,i)\}_{i=1}^{a}$ and $\{\bm{\mathrm{nv}}_{2}(k,i)\}_{i=1}^{b}$, where $k$ is the layer index and $a\geq b$, i.e., one model has more neurons than the other. We concatenate the neuron vectors of the two models into two matrices $\bm{H}_{1}\in\mathbb{R}^{n\times a}$ and $\bm{H}_{2}\in\mathbb{R}^{n\times b}$, where $n$ is the number of samples. The main obstacle here is that we cannot directly compare the neuron vectors in $\bm{H}_{1}$ and $\bm{H}_{2}$ because there is no apparent functionality correspondence between the neurons of the two models with different structures. To tackle this problem, we construct a linear transformation $\bm{P}\in\mathbb{R}^{b\times a}$ to project neuron vectors in $\bm{H}_{1}$ into a space of the same dimension as $\bm{H}_{2}$ so that the metrics we designed above become applicable to $\bm{H}_{1}{\bm{P}}^{\intercal}$ and $\bm{H}_{2}$. The linear transformation $\bm{P}$ should preserve the difference among independently trained models in terms of neuron functionality, as well as the similarity between victim models and surrogate models, to enable reuse relation detection between heterogeneous models.
@@
where $\lVert\cdot\rVert_{F}$ is the Frobenius norm. Intuitively, neuron vectors of reference models are largely linearly independent of victim model’s neuron vectors since they are independently trained, so the distances between victim models and reference models will still remain relatively large after the linear transformation.
@@
$\bm{p}_{i}=\operatorname*{argmin}_{\bm{w}_{i}}\lVert\bm{H}_{1}\bm{w}_{i}-\bm{\mathrm{nv}}_{2}(k,i)\rVert_{2},$ (10)
@@ IV Benchmark
In order to evaluate our approach comprehensively, we create a benchmark named Reuse Zoo consisting of 250 models for model reuse detection. This benchmark contains 4 pre-trained models (PTMs), 126 surrogate models, and 120 reference models with four popular classification model architectures, including ResNet-18 [35], ResNet-50 [35], MobileNet V2 [36], and GoogLeNet [37].
@@
Knowledge distillation and model extraction are typically used to train smaller surrogates with different architectures from the source model. Since the capacity of ResNet-50 is significantly larger than that of the other three architectures, we train three knowledge distillation surrogates and three model extraction surrogates (with ResNet-18, MobileNet V2, and GoogLeNet architectures) for each ResNet-50 transfer learning model. We use the vanilla method [33] for knowledge distillation and the Knockoff method [9] for model extraction. These models constitute the knowledge distillation surrogate set with 9 models and the model extraction surrogate set with 9 models. The pruning, quantization, fine-tuning, retraining, knowledge distillation, and model extraction surrogate sets contain 114 surrogate models whose source victim models are the transfer learning surrogate models. Together with the transfer learning surrogate set, we obtain a total of 126 surrogate models using these seven model reuse techniques.
We construct 120 reference models in the following way. We independently train ten models from scratch using different weight initialization for each model architecture (four architectures) and each dataset (three datasets). All surrogate and reference models are well-trained and have qualified performance. The complete list of models in the Reuse Zoo benchmark is shown in Table I, where several examples of the models are also given.
This benchmark covers four popular model architectures, ranging from small networks like MobileNet V2 to large networks like ResNet-50, and various model reuse techniques, notably including heterogeneous reuse methods that can change the model structure. We note that the authors of ModelDiff [24] have proposed a similar DNN reuse detection benchmark referred to as ModelReuse. Compared with ModelReuse, Reuse Zoo (1) is constructed with three commonly used classification datasets while ModelReuse uses two fine-grained image classification (FGIC) datasets that are less representative; (2) includes two more model architectures than ModelReuse, namely ResNet-50 and GoogLeNet, where ResNet-50 is a significantly larger model; (3) covers more reuse techniques, such as fine-tuning and retraining on different layers, which are absent from ModelReuse; (4) is larger, more balanced (126 surrogate models and 120 reference models vs. 84 and 28), and of higher quality (5% accuracy gap between homogeneous models vs. 25% in ModelReuse).
V Empirical Evaluation
Parameter setting. We implement NFARD using PyTorch [42] and release its source code along with the benchmark. [^4] Artifacts available at https://figshare.com/s/691d5af98dc715cf7009 NFARD has four parameters: the test suite size $n$, the user-specified parameter $\alpha$ for decision criteria, the distance metrics weights $\{w_{i}\}$, and the choice of target layer under the white-box setting. If not specified below, we use a test suite size of $n=1000$, and we empirically set the weights of Euclidean and cosine distances as 1 and 120, respectively. $\alpha$ is set to 0.85 for the black-box case and 3.5 for the white-box case. The target layer is set to the layer at about 25% depth. We will discuss the choice of $\alpha$ and target layer later.
V-A RQ1: Performance of NFARD
V-A1 Benchmark evaluation result
@@
Reuse type Detected Decision $\mathrm{BED}$ $\mathrm{BCD}$ Detected Decision $\mathrm{ED}^{(k)}$ $\mathrm{CD}^{(k)}$
@@ V-A2 Comparison with baselines
We also present the evaluation results of NFARD and baseline methods on the ModelReuse benchmark in Table V. NFARD achieves a higher F1-score compared to the baseline methods, particularly excelling in heterogeneous reuse cases. However, the reference models in the ModelReuse benchmark exhibit lower accuracy than the victim models (e.g., 55% vs. 80%), making them less suitable for computing reference threshold values for decision-making due to their reduced similarity to the victim models. As a result, all methods evaluated on the ModelReuse benchmark, including NFARD, show a lower F1-score and a higher false positive rate compared to their performance on the Reuse Zoo benchmark. Overall, the results highlight the effectiveness of NFARD in detecting model reuse relations and underscore the comprehensiveness of the Reuse Zoo benchmark relative to the ModelReuse benchmark.
@@
TABLE V: Comparison of NFARD with baseline on the ModelReuse benchmark. The best F1-score is highlighted.
Reuse type NFARD IPGuard ModelDiff DeepJudge
Pruning - 0.2 12/12 10/12 12/12 12/12
Pruning - 0.5 12/12 9/12 12/12 12/12
Pruning - 0.8 6/12 2/12 9/12 12/12
Quantization - qint8 12/12 12/12 12/12 12/12
Transfer learning 9/12 0/12 10/12 0/12
Knowledge distillation 10/12 0/12 9/12 12/12
Model stealing 12/12 0/12 2/12 0/12
Positive 73/84 33/84 66/84 60/84
Negative 3/28 0/28 4/28 0/28
F1-score 0.913 0.564 0.857 0.833
@@ V-B RQ2: Application Scope
TABLE VI: Qualitative comparison of the application scope of NFARD with adversarial example-based methods.
@@
TABLE VII: Test suite generation time (in seconds) of reuse detectors. The shortest generation time is highlighted.
@@
We have shown that NFARD is capable of detecting more reuse types than other DNN reuse detection methods by employing the linear transformation method. To answer the second research question, we further compare the application scope of NFARD with other methods in the following aspects: model access, scalability, and label information. We provide a brief summary in Table VI and describe them in detail.
@@ V-B2 Scalability
Another important factor that affects the application scope is scalability, which is mainly limited by the efficiency of test suite generation as the metric computation and decision-making process usually takes no more than a few seconds. We report the average test suite generation time for different model architectures of NFARD and other adversarial example-based methods (with $n=1000$) in Table VII. The test suite generation process for NFARD is the same in black-box and white-box cases, while DeepJudge takes two different procedures for test suite generation in each case. NFARD is $2\sim 99$ times faster than the baseline reuse detectors, and its efficiency is also less influenced by the complexity of the model. Since NFARD directly selects samples from the training dataset, the test suite generation time stays almost the same as the number of test cases grows. On the contrary, the test suite generation time of adversarial example-based reuse detectors grows linearly with the number of test cases.
@@ V-C RQ3: Influencing Factors
Next, we explore how different configurations affect the performance of NFARD. First, we perform an ablation study to evaluate the effectiveness of the logarithm approximation, as well as the necessity of the distance metrics. Then, we test the performance with different parameters, including decision parameter $\alpha$, target layer choice, and test suite size $n$.
V-C1 Logarithm approximation and distance metrics
TABLE VIII: Evaluation results of NFARD without taking the logarithm under the black-box setting.
Reuse type Detected Decision $\mathrm{BED}$ $\mathrm{BCD}$
@@ V-C1 Logarithm approximation and distance metrics
We report the evaluation results of NFARD without taking the logarithm in the black-box setting in Table VIII. There is a significant drop in performance without the logarithm approximation, especially for knowledge distillation and model extraction reuse cases. The decision value also decreases for most reuse cases, suggesting that the logarithm approximation method does amplify the difference between independently trained models and surrogate models in the black-box case. The F1-score of NFARD using only the BED metric or the BCD metric under the black-box setting is 0.9756 and 0.9520, respectively, which is lower than the F1-score of 0.984 using both metrics. This indicates that the distance metrics are complementary to each other and contribute to the overall performance.